### v1

# Understanding Action Effects through Instrumental Empowerment in Multi-Agent Reinforcement Learning

Abstract
To reliably deploy Multi-Agent Reinforcement Learning (MARL) systems, it is crucial to understand individual agent behaviors within a team. While prior work typically evaluates overall team performance based on explicit reward signals or learned value functions, it is unclear how to infer agent contributions in the absence of any value feedback. In this work, we investigate whether meaningful insights into agent behaviors can be extracted that are consistent with the underlying value functions, solely by analyzing the policy distribution. Inspired by the phenomenon that intelligent agents tend to pursue convergent instrumental values, which generally increase the likelihood of task success, we introduce Intended Cooperation Values (ICVs), a method based on information-theoretic Shapley values for quantifying each agent’s causal influence on their co-players’ instrumental empowerment. Specifically, ICVs measure an agent’s action effect on its teammates’ policies by assessing their decision uncertainty and preference alignment. The analysis across cooperative and competitive MARL environments reveals the extent to which agents adopt similar or diverse strategies. By comparing action effects between policies and value functions, our method identifies which agent behaviors are beneficial to team success, either by fostering deterministic decisions or by preserving flexibility for future action choices. Our proposed method offers novel insights into cooperation dynamics and enhances explainability in MARL systems.
[…]
3982
1 Introduction
The rise of Multi-Agent Reinforcement Learning (MARL) has led to many real-world problems being reformulated as games. MARL offers the greatest benefits when combined with powerful function approximators, such as deep neural networks. Beyond robust performance, it is crucial to understand these models by extensively testing and analyzing their behavior in simulation environments to ensure safe deployment in real-world systems. This need has driven attention toward post-hoc explanation methods, which treat models as black boxes without constraining their complexity. A game-theoretic method in this category is the Shapley Value (SV), which was originally designed to fairly distribute the overall payoff among players in a cooperative game [6]. It has since been adapted as a feature attribution technique to explain why specific predictions are made in a deep learning model [36].
In MARL, the reward function is fundamental in shaping agent behavior, with Schmidhuber’s artificial curiosity [33] inspiring the design of intrinsic rewards to foster cooperation. As the majority of MARL algorithms (see, e.g., [30]) rely on learned value functions to estimate expected rewards in a given state, the SV naturally emerges as a credit assignment method in fully cooperative MARL scenarios. The credit assignment problem in MARL concerns identifying each agent’s contribution to the team’s overall performance. Conversely, in the absence of both value functions and reward signals, a critical question arises of how, and more importantly which, credit should be allocated, irrespective of the game type (cooperative or competitive). In this setting of missing explicit value feedback, we investigate whether SVs can be adapted to extract meaningful information about agent behavior on performance, i.e., information that remains consistent with the underlying value function, while assuming access only to the policy function.
To identify a meaningful candidate for credit, we take inspiration from the famous statement by von Foerster [41]: "I shall act always so as to increase the total number of choices", thereby advocating for individual empowerment. In a related vein, Omohundro [28] derived the fundamental tendency of optimal agents to seek power. Both philosophies align with the instrumental convergence thesis by Bostrom [5], which posits that several instrumental values are universally beneficial to increase task success and thus likely to be pursued by many intelligent agents. It has been theoretically [4, 39] and empirically [10, 20] shown that access to and control over a wide range of future states (or options) serves as a proxy for preparedness. This makes it a convergently instrumental objective, as formally defined by the empowerment function [16] or by the policy entropy [14]. Given the evidence that state space exploration is generally beneficial during training, we question whether the same holds true during inference in multi-agent settings. Concretely, is providing agents with more options truly beneficial for task success?
updownleftright011
updownleftright011
Figure 1: Left: a toy environment where both players must reach the target (star) to earn a shared reward, while gray cells block the path. Blue and red arrows indicate potential moves (choices). Right: probabilities (y-axis) over actions (x-axis) of P2 before (top) and after (bottom) P1 steps on the key cell and opens the green lock. By opening the green key, P2 becomes more certain about choosing the leftward action, whereas unlocking the gray locks would provide additional movement options that are clearly not beneficial for P2.
Drawing on the classic exploration-exploitation dilemma in Reinforcement Learning (RL) [37], we argue that exploiting certain situations with high certainty may be more instrumentally valuable than preserving many choices. We illustrate this concept in Fig. 1. The maze on the left shows a toy environment where two agents must reach a target (star) to receive a shared reward of one. The histograms on the right depict the probability distributions over movements actions for P2 before (top) and after (bottom) P1 acted. Consider the case where stepping on the key opens the green lock only. In the current state, the lock cell to the left of P2 is closed, which makes P2 more likely to move downward, while still assigning some probability to the leftward action in anticipation that P1 might open the lock. Once P1 steps on the key cell, this action causally affects P2’s action space by unlocking the left path. As a result, P2 not only gets more movement options, but also becomes more certain in exploiting the open path by assigning more weight toward the left action. This highlights how P2’s (increased) success is directly tied to P1’s altruistic behavior, which would correspond to an increased value under an optimal value function. In contrast, if the key were to unlock the gray locks, it would provide additional movement options that are clearly not beneficial for achieving its goal. Even if the game involves simultaneous decision-making and P2 chooses to go down, P1 would merit credit for his intention to open the path and contribute to P2’s success. In addition to analyzing effects on decision uncertainty, we examine the impact on preference alignment between agents. To do so, we perform a state exchange intervention by placing P1 in P2’s position and v.v. to assess whether agents exhibit coordinated consensus through mutual anticipation of similar behavior or whether they favor divergent strategies. Credit is attributed to actions that enhance alignment and promote consensus.
Motivated by this, we introduce a causal action attribution method called Intended Cooperation Value (ICV), which quantifies how cooperatively an agent behaves by measuring the causal effect of its actions on the instrumental empowerment of its teammates, using only access to their policies. By adapting causal, information-theoretic Shapley Values, ICV uses the entropy of the policy distributions to assign credit to an agent based on its intended contribution to reducing decision uncertainty and fostering preference alignment among co-players. Empirical evaluations across cooperative, competitive, and mixed-motive settings demonstrate that ICV reliably attributes credit to behaviors that increase the likelihood of task success and aligns with standard value-based SVs. Our approach provides a deeper understanding of agent interactions, particularly in scenarios where success is not explicitly measurable.
2 Background
Let $(\Omega,\mathcal{F},\Pr)$ be a probability space. Let $X:\Omega\rightarrow\mathcal{X}$ be a discrete Random Variable (RV) taking value $x$ in a set $\mathcal{X}$, where $p(x)=\Pr(X=x)$ is the Probability Mass Function (PMF) and $\mathcal{P}(\mathcal{X})$ is the set of probability distributions over $\mathcal{X}$. We restrict our attention to finite sets $\mathcal{X}$. The entropy of $X$ is given by $H(X)=-\sum_{x\in\mathcal{X}}p(x)\log p(x)$. For two PMFs $p$ and $q$, the Kullback–Leibler divergence is $D_{KL}(p||q)=\sum_{x\in\mathcal{X}}p(x)\log\frac{p(x)}{q(x)}$.
2.1 Reinforcement Learning
The interactions among $n$ concurrently acting agents are typically modeled using either a Markov Game (MG) [22] or an Extensive-Form Game (EFG) [29]. While in EFGs actions are selected and executed sequentially, in MGs actions are selected simultaneously but their execution may be either parallel or sequential, depending on whether the environment supports parallelization. If not, actions are executed based on a predefined or randomized agent ordering.
Assumption 1.
Without loss of generality, we focus on MGs, as simultaneous decision-making is commonly used in MARL systems [1].
Assumption 1 implies that we consider domains with full state observability, though our approach can trivially be extended to partial observability and sequential decisions.
Definition 1.
An MG is defined by the tuple
$\mathcal{G}_{m}\coloneqq(\mathcal{N},\mathcal{S},\{\mathcal{A}^{i}\}_{i=1}^{n},P,\rho,\{R^{i}\}_{i=1}^{n},\gamma),$
where $\mathcal{N}=\{1,\dots,n\}$, with $n\in\mathbb{N}$, is the set of agents, $\mathcal{S}$ is the set of admissible states, $\mathcal{A}^{i}$ is the set of actions available to agent $i\in\mathcal{N}$, with $\boldsymbol{\mathcal{A}}=\prod_{i=1}^{n}\mathcal{A}^{i}$ being the joint action set, $P:\mathcal{S}\times\boldsymbol{\mathcal{A}}\to\mathcal{P}(\mathcal{S})$ is the state transition probability kernel, $\rho\in\mathcal{P}(\mathcal{S})$ is the initial state distribution, $R^{i}:\mathcal{S}\times\boldsymbol{\mathcal{A}}\times\mathcal{S}\to\mathbb{R}$ is the reward function of agent $i$, and $\gamma\in[0,1)$ is the discount factor controlling the effect of future rewards. At each time step $t\in\mathbb{N}$, let the RV $S_{t}$ describe the state of the Markov game, and the RV $A_{t}^{i}$ denote the action taken by agent $i$. Starting from an initial state $s_{0}\sim\rho$, all agents observe the current state $s_{t}\in\mathcal{S}$ and simultaneously select an action $a_{t}^{i}\in\mathcal{A}^{i}$ according to their individual policies $\pi^{i}(a\mid s_{t})$, forming a joint action $\boldsymbol{a}_{t}=(a_{t}^{1},\dots,a_{t}^{n})\in\boldsymbol{\mathcal{A}}$ drawn from the joint policy $\boldsymbol{\pi}(\boldsymbol{a}\mid s_{t})=\prod_{i=1}^{n}\pi^{i}(a\mid s_{t})$. The MG then transitions to a new state $s_{t+1}$ according to $P(s_{t+1}\mid s_{t},\boldsymbol{a}_{t})$, and each agent $i$ receives a reward $R^{i}(s_{t},\boldsymbol{a}_{t},s_{t+1})$. The performance of each agent $i$ can be measured by means of a value function defined as:
$V^{i}(s)\coloneqq\mathbb{E}_{\boldsymbol{\pi},P}\left[\sum_{t=0}^{\infty}\gamma^{t}R^{i}(S_{t},\boldsymbol{A}_{t},S_{t+1})\mid S_{0}=s\right].$ (1)
We restrict our attention to environments satisfying the MG dynamics as per Def. 1. Henceforth we, use the terms environment, game, and MG interchangeably.
Definition 2.
Multi-agent systems (game types) can be categorized as [32]: i) Cooperative: all agents $i\in\mathcal{N}$ share a global reward $R=R^{i}$. ii) Competitive: $\sum_{i=1}^{n}R^{i}=0$ for any state transition (zero-sum MG). iii) Mixed-motive: having individual rewards that induce cooperative and competitive motivations.
Assumption 2.
We consider pre-trained Actor-Critic [17] models under the MG framework using Centralized Training with Decentralized Execution (CTDE) [24], the state-of-the-art paradigm for MARL. Further, we consider stochastic policies, which are generally more robust and tend to perform better in real-world-like environments compared to deterministic policies [46].
From Definition 2, we emphasize the flexibility of MGs to represent all game types [9]. Assumption 2 implies that agents act with individual policies (actors), and in fully cooperative settings, they share a global value function $V(s)$, also called centralized critic [2]. We use $\boldsymbol{-i}$ to denote the set of all agent indices excluding agent $i$.
2.2 Shapley Value
We build on prior work applying SVs for feature attribution in single-agent RL and Machine Learning (ML) models. Therefore, let $Z$ represent the state in RL or the input in ML.
Definition 3.
Let $\mathcal{M}=\{1,\dots,m\}$ be the set of features and $\mathcal{Z}=\prod_{i=1}^{m}\mathcal{Z}_{i}$ the input space, so that any input $z$ can be represented as an ordered set $z=\{z_{i}\mid z_{i}\in\mathcal{Z}_{i}\}_{i=1}^{m}$, with $p(z)$ being the data distribution. Then, any subset $\mathcal{K}\subseteq\mathcal{M}$ yields a partial observation $z_{\mathcal{K}}=\{z_{i}\mid i\in\mathcal{K}\}$, where $\bar{\mathcal{K}}=\mathcal{M}\setminus\mathcal{K}$ is its complement.
Definition 4.
The Shapley Value [34] is the unique method that satisfies the desirable fairness properties of efficiency, symmetry, additivity, and null player (see Appendix B). It allocates credit by computing the average marginal contribution $\Delta v(\cdot)$ of player (or feature) $i$ to the outcome of a cooperative game (or prediction) over all coalitions $\mathcal{K}\subseteq\mathcal{M}$ using a characteristic function $v:2^{\mathcal{M}}\rightarrow\mathbb{R}$, as follows:
$\phi_{i}(v)=\sum_{\mathcal{K}\subseteq\mathcal{M}\setminus\{i\}}\frac{|\mathcal{K}|!\,(|\mathcal{M}|-|\mathcal{K}|-1)!}{|\mathcal{M}|!}\underbrace{\left[v(\mathcal{K}\cup\{i\})-v(\mathcal{K})\right]}_{\Delta v_{i}(\cdot)}\,.$
3 Related Work
There has been considerable work on understanding and explaining RL, primarily aimed at answering why specific actions were taken. Existing approaches include replacing the policy network with transparent models, such as decision trees [38], or incorporating auxiliary reasoning mechanisms like counterfactual analysis [45], structural causal models [26], or interpretable reward components [23]. The SV has been used as a feature attribution method in Machine Learning (ML) [12, 47], in single-agent RL [3, 35], and in MARL to assess agent contributions based on rewards [27]. Beyond explanatory purposes, SVs have also been adapted as training mechanisms in MARL, mainly to address the credit assignment problem [7, 43, 42]. In parallel, various information-theoretic intrinsic rewards have been proposed to foster multi-agent cooperation [8, 25].
So far, no work has examined whether causal influences on agent policies, measured via information-theoretic SVs, correspond to changes in the value function. We are the first to explore if analyzing solely the policy function provides meaningful insights into an agent’s behavior and its impact on co-players’ decisions, and if these insights align with overall team performance.
4 Method
We aim for a method applicable during inference without requiring a separately trained model or modifications to the underlying MARL architecture. We introduce our approach as follows: i) a framework for modeling simultaneous action selection with sequential action execution (Sec. 4.1), ii) an adaptation of SVs to isolate an agent’s causal effect and measure its marginal contribution during gameplay (Sec. 4.2), iii) characteristic functions for quantifying the instrumental value of an agent’s action on its co-players (Sec. 4.3).
4.1 Game Model
The ordinary MG framework makes it hard to infer individual action effects on others. For instance, in Fig. 1, imagine an agent P3 is positioned below the green lock instead of the wall and P3 could also step onto the lock cell when it gets opened by P1. By analyzing only $S_{t}$ and $S_{t+1}$, one cannot determine individual contributions of P1 and P3 on P2’s decision-making. To overcome this limitation, we propose a new game model, the Sequential Value Markov Game (SVMG).
Definition 5.
An SVMG is defined by the tuple
$\mathcal{G}_{s}\coloneqq(\mathcal{N},\mathcal{S},\{\mathcal{A}^{i}\}_{i=1}^{n},\{P^{i}\}_{i=1}^{n},\mathcal{D},\rho),$
where $\mathcal{N},\mathcal{S},\{\mathcal{A}^{i}\}_{i=1}^{n}$, and $\rho$ are the same variables as for $\mathcal{G}_{m}$ in Def. 1. A transition kernel $P^{i}:\mathcal{S}\times\mathcal{A}^{i}\rightarrow\mathcal{P}(\mathcal{S})$ considers only one agent action. $\mathcal{D}\in\mathcal{P}(\Sigma)$ specifies a distribution over the set of all $n!$ permutations $\Sigma\coloneqq\{\sigma:\mathcal{N}\hookrightarrow\mathcal{N}\}$.
Consider Fig. 2 to better elaborate on the SVMG framework. Without loss of generality we assume that $\mathcal{S}=\prod_{i=1}^{n}\mathcal{S}_{i}$ so that any state RV $S:\Omega\rightarrow\mathcal{S}$ can be decomposed into states of individual agents $S=(S^{1},S^{2},\ldots,S^{n})$, where $S^{i}:\Omega\rightarrow\mathcal{S}^{i}$. Then, actions are executed sequentially in $\mathcal{G}_{s}$ according to an ordering $\sigma\sim\mathcal{D}$, where $\mathcal{D}$ is either uniform or deterministic. While $i\in\mathcal{N}$ is an agent index, let $k\in\mathcal{N}$ be a sub-step index iterating over $\sigma$, where $\sigma(k)$ denotes the agent assigned to the $k$-th position in the ordering. Thus, agent $i$ acts at sub-step $k$ if and only if $i=\sigma(k)$.
Definition 6.
We define intermediate state RVs $S_{t,(k)}$ and $S_{t}$ for $t$ and $k$ in the following way. Let $S_{0}=S_{0,(0)}\sim\rho$. For each $t$ we construct $S_{t,(k)}$ such that $S_{t,(k)}|S_{t,(k-1)},A_{t}^{\sigma(k)}\sim P^{\sigma(k)}$, and $S_{t+1}=S_{t,(n)}$. Each kernel $P^{i}$ has a property that it updates only state of player $i$ and the states of other agents stays fixed, i.e., $P^{i}(s_{k}|s_{k-1},a^{i})=0$ for all $j\neq i$ when $s_{k}^{j}\neq s_{k-1}^{j}$. For each $t$ a new $\sigma$ is sampled iidiid from $\mathcal{D}$.
By Definition 6, agent $i$ got its local state $S_{t,(k)}^{i}$ is updated at sub-step $k$ if it belongs to the first $k$ actors, i.e.,
$S_{t,(k)}^{i}=\begin{cases}S_{t+1}^{i},&\sigma^{-1}(i)\leq k\\ S_{t}^{i},&\sigma^{-1}(i)>k\end{cases}$ (2)
Assumption 3.
Def. 6 implies that we assume $\mathcal{G}_{s}$ to be locally Markovian, i.e., for each $i$ and step $k$, the kernel $P^{i}$ depends only on the immediately preceding state $s_{t,(k-1)}$ and the action $a_{t}^{i}$:
$\Pr(S_{t,(k)}=s_{t,(k)}\mid\{S_{t,(j)}=s_{t,(j)}\}_{j<k},\{A_{t}^{j}=a_{t}^{j}\}_{j=1}^{n},\sigma(k)=i)$
$\eqqcolon P^{i}(s_{t,(k)}\mid s_{t,(k-1)},a_{t}^{i}).$
$S_{t}$$S_{t+1}$$\boldsymbol{A}_{t}$$S_{t,(0)}$$S_{t,(1)}$$S_{t,(2)}$$S_{t,(3)}$$A_{t}^{\sigma(1)}$$A_{t}^{\sigma(2)}$$A_{t}^{\sigma(3)}$$S^{\sigma(1)}_{t}$$S^{\sigma(2)}_{t}$$S^{\sigma(3)}_{t}$$S^{\sigma(1)}_{t+1}$$S^{\sigma(2)}_{t}$$S^{\sigma(3)}_{t}$$S^{\sigma(1)}_{t+1}$$S^{\sigma(2)}_{t+1}$$S^{\sigma(3)}_{t}$$S^{\sigma(1)}_{t+1}$$S^{\sigma(2)}_{t+1}$$S^{\sigma(3)}_{t+1}$
Figure 2: Transformation of MG (upper-level, saturated colors) to SVMG (low-level, faded colors) for $n=3$ agents, following an execution order $\sigma$. Red lines between states mean correspondences, while dashed lines mean decomposition into individual components.
Assumption 4.
We assume the underlying MG can be expressed by the SVMG by fulfilling the following conditions:
  1. i)
    The environment possess an internal ordering of agents, according to which actions are executed sequentially.
  2. ii)
    The joint kernel $P$ of $\mathcal{G}_{m}$ is decomposable by the per sub-step kernels $P^{\sigma(k)}$ of $\mathcal{G}_{s}$, such that by marginalizing over all orderings $\sigma\in\Sigma$ and intermediate states $\mathbf{s}=(s_{t,(1)},\dots,s_{t,(n-1)})$, we recover the joint kernel satisfying $s_{t,(0)}=s_{t}$ and $s_{t,(n)}=s_{t+1}$:
    $P(s_{t+1}\mid s_{t},\boldsymbol{a}_{t})=\mathbb{E}_{\sigma}\sum_{\mathbf{s}\in\mathcal{S}}\prod_{k=1}^{n}P^{\sigma(k)}\left(s_{t,(k)}\mid s_{t,(k-1)},a_{t}^{\sigma(k)}\right)$ (3)
While Assumption 4 may hold only in simple environments, it enables our framework to be applied online during execution. In such cases, intermediate states $S_{t,(k)}$ can be accessed while actions are being processed. In contrast, for more complex games that support parallel execution and thereby violate Assumption 4, our SVMG framework remains applicable offline. By storing the global state trajectory $\{S_{t}\}_{t=0}^{T}$ during inference, the intermediate sequences $\{S_{t,(k)}\}_{k=1}^{n}$ can be reconstructed by Eq. (2) in a post-processing step with $\sigma$ injected externally.
Causality.
$\mathcal{G}_{s}$ in Fig. 2 represent causal Directed Acyclic Graphs (DAGs), where edges like $S_{t,(k)}\rightarrow A^{\sigma(k+1)}_{t}$ indicate direct causal relationships. However, for $k>1$, $S_{t,(k)}$ may not directly cause $A^{\sigma(k+1)}_{t}$, since joint actions $\boldsymbol{A}_{t}$ are pre-determined at $S_{t}=S_{t,(0)}$, and agents do not react to all intermediate states. To account for this, we perform an intervention [31] by fixing $A^{\sigma(k)}_{t}$ to $a^{\sigma(k)}_{t}$ as $\mathrm{do}(A^{\sigma(k)}_{t}=a^{\sigma(k)}_{t})$, independently of its parents. Graphically, this corresponds to removing incoming arcs to $A^{\sigma(k)}_{t}$, yielding a surgery marked by grayish arcs in Fig. 2. Viewing $\mathcal{G}_{s}$ as a causal DAG enables post-interventional queries to analyze local inter-agent influences, as similarly done by Du et al. [8].
4.2 Action Attribution
We present our adaptation of ordinary SVs to isolate an agent’s causal effect and draw inspiration from Causal Shapley Values, which have been applied to ML predictions. Based on Def. 3 and 4, Heskes et al. [15] modify the data distribution by intervening on coalition features and propose $v=p(z_{\bar{\mathcal{K}}}\mid\text{do}(Z_{\mathcal{K}}=z_{\mathcal{K}}))$. In contrast, Frye et al. [11] assign nonzero weight $w$ only to causal orderings of $z_{\mathcal{K}}$, allowing feature importance to be distributed selectively. In contrast to $v$ from Def. 4, which typically measures the effects on the overall game outcome, we aim for measuring effects of players during the game. We extend $v:\Sigma\rightarrow\mathbb{R}$ to consider the current game situation and propose the characteristic function $\nu:\mathcal{S}\times\Sigma\rightarrow\mathbb{R}$, for which concrete definitions are provided in Sec. 4.3.
Definition 7.
Given an order $\sigma$ and $j\succ_{\sigma}i$ meaning $j$ succeeds $i$ in $\sigma$, we define agent specific coalitions as $\mathcal{C}_{\sigma}^{i}\coloneqq\{j:j\succ_{\sigma}i\}.$
Definition 8.
Let agent $i=\sigma(k)$ be the acting player. Given an order $\sigma\in\Sigma^{i}_{+}$, where $\Sigma^{i}_{+}=\{\sigma\in\Sigma:\mathcal{C}_{\sigma}^{i}\neq\emptyset\}$, we define the marginal contribution of agent $i$ as the effect on members of coalition $\mathcal{C}_{\sigma}^{i}$ during action processing steps k1k-1 and $k$, as:
$\Delta\nu\left(\mathcal{C}_{\sigma}^{i},s_{t,(k)}\right)\coloneqq\nu\left(\mathcal{C}_{\sigma}^{i},s_{t,(k)}\right)-\nu\left(\mathcal{C}_{\sigma}^{i},s_{t,(k-1)}\right).$
In Def. 8, the inclusion of agent $i$ is not explicit as in $\Delta v(\cdot)$ from Def. 4, i.e., $\mathcal{C}_{\sigma}^{i}\cup\{i\}$, but occurs through executing its action $a_{t}^{i}$ and transitioning to $s_{t,(k)}\sim P^{i}(\cdot\mid s_{t,(k-1)},\text{do}(A_{t}^{i}=a_{t}^{i}))$ according to Def. 6. This serves two purposes: i) preserving natural game execution while measuring an agent’s effect within the causal chain of $\mathcal{G}_{s}$, ii) capturing agent $i$’s intention on succeeding agents in $\mathcal{C}_{\sigma}^{i}$ whose actions have not yet been processed to not attribute their impact.
This is visualized in Fig. 3, where the affected states $\{S_{t,(\cdot)}^{j}\}_{j\in\mathcal{C}_{\sigma}^{i}}$ used for computing $\Delta\nu\left(\mathcal{C}_{\sigma}^{i},s_{t,(k)}\right)$ match the processed action $a_{t}^{i}$. Accordingly, the only ordering without causal effect is the empty coalition $\mathcal{C}_{\sigma}^{i}=\emptyset$, which is explicitly excluded in Def. 8 by $\Sigma_{+}^{i}$. To quantify an agent’s average effect during the game, the weight for each $\sigma\in\Sigma_{+}^{i}$ is given as $w\coloneqq\mathrm{Pr}(\sigma)=\frac{1}{\lvert\Sigma^{i}_{+}\rvert}=\frac{1}{n!-(n-1)!}$.
Definition 9.
We define the Intended Cooperation Value (ICV) to assign credit to agent $i$ based on its average action effect on the members of coalition $\mathcal{C}_{\sigma}^{i}$ during the time horizon $T$ as:
$\Phi_{i}\left(\nu\right)\coloneqq\frac{1}{T}\sum_{t=0}^{T-1}\sum_{\sigma\in\Sigma_{+}^{i}}w\cdot\Delta\nu\left(\mathcal{C}_{\sigma}^{i},s_{t}\right).$
In practice, it is computationally infeasible to consider all coalitions for large $n$ [40]. As in previous works [7], we use Monte Carlo sampling to get an unbiased estimator $\hat{\Phi}$ by sampling a new order $\sigma_{t}$ for each step $t$ and by assuming $T\gg n$, as follows:
$\hat{\Phi}_{i}\left(\nu\right)\coloneqq\frac{1}{T}\sum_{t=0}^{T-1}\cdot\Delta\nu\left(\mathcal{C}_{\sigma_{t}}^{i},s_{t}\right),\quad\sigma_{t}\sim\mathcal{U}_{\Sigma_{+}^{i}}.$ (4)
$S^{\sigma(1)}_{t,(0)}$$S^{\sigma(2)}_{t,(0)}$$S^{\sigma(3)}_{t,(0)}$$a^{\sigma(1)}_{t}$$S^{\sigma(1)}_{t,(1)}$$S^{\sigma(2)}_{t,(1)}$$S^{\sigma(3)}_{t,(1)}$$a_{t}^{\sigma(2)}$$S^{\sigma(1)}_{t,(2)}$$S^{\sigma(2)}_{t,(2)}$$S^{\sigma(3)}_{t,(2)}$$a_{t}^{\sigma(3)}$$S^{\sigma(1)}_{t,(3)}$$S^{\sigma(2)}_{t,(3)}$$S^{\sigma(3)}_{t,(3)}$
Figure 3: Intermediate states $S_{t,(\cdot)}^{\sigma(\cdot)}$ matching the color of the processed action $a_{t}^{\sigma(k)}$ are used for computing the marginal contribution $\Delta\nu$. Notice that $S_{t,(1)}^{\sigma(3)}$ is used for the effects of both $a_{t}^{\sigma(1)}$ and $a_{t}^{\sigma(2)}$. Arcs entering $a_{t}^{\sigma(k)}$ are explicitly ignored due to intervention.
4.3 Characteristic Functions
We present various definitions for $\nu$ based on Def. 3 and 4.
Value-based.
Beechey et al. [3] derived the correct application of SVs in single-agent RL by introducing the following characteristic functions, with $p^{\pi}(z^{\prime}\mid z_{\mathcal{K}})$ as the state occupancy distribution:
$v_{V}(\mathcal{K})=\sum_{z^{\prime}\in\mathcal{Z}}p^{\pi}(z^{\prime}\mid z_{\mathcal{K}})V(z^{\prime}),$ (5)
$v_{\pi}(\mathcal{K})=\sum_{z^{\prime}\in\mathcal{Z}}p^{\pi}(z^{\prime}\mid z_{\mathcal{K}})\pi(a\mid z^{\prime}).$ (6)
They showed that SVs with $v_{V}(\mathcal{K})$ capture the contribution of state features to the value function (Eq. (1)) as a static predictor, and with $v_{\pi}(\mathcal{K})$, their influence on the probability of selecting action $a$. In contrast, $\nu(\mathcal{C}_{\sigma}^{i},s_{t,(k)})$ does not evaluate subsets of features, but full realizations of intermediate states $s_{t,(k)}$, where feature contributions emerge from the effect of an agent’s actions on the state.
Definition 10.
The characteristic value-based function is defined as:
$\nu_{v}(\mathcal{C}_{\sigma}^{i},s_{t,(k)})\coloneqq\sum_{j\in\mathcal{C}_{\sigma}^{i}}V^{j}(s_{t,(k)}),$
which reduces to $\nu_{v}(\mathcal{C}_{\sigma}^{i},s_{t,(k)})=V(s_{t,(k)})$ in the case of centralized critics, with $V^{j}(\cdot)$ being defined in Eq. (1).
Entropy-based.
As noted in the introduction, access to many future states is instrumentally valuable. The size of future accessible states can be quantified by $\lvert\mathcal{S}_{+}\rvert$, where $\mathcal{S}_{+}=\{s_{t+1}\in\mathcal{S}:P(s_{t+1}\mid s_{t})>0\}$. Based on this, Franzmeyer et al. [10] empirically validated the use of $H(A_{t}\mid s_{t})$ as a proxy for $\lvert\mathcal{S}_{+}\rvert$ to capture an agent’s future choices. Similarly, Watson et al. [44] introduced information-theoretic SVs via $v=H(Y\mid z_{\mathcal{K}})$ to measure predictive uncertainty in ML, with $Y$ being the output. Building on both, we argue that in some multi-agent settings, it may be more valuable for an agent to be certain about what to do, i.e., committed to a specific strategy. Since $H(A_{t,(k+1)}^{i}\mid s_{t,(k)})\leq\log|\mathcal{A}^{i}|$ [13] quantifies the spread of $\pi^{i}(\cdot\mid s_{t,(k)})$, we define its complement to this bound as a measure of peakedness: $\mathrm{H}(A_{t,(k+1)}^{j})\coloneqq\log\lvert\mathcal{A}^{j}\rvert-H(A_{t,(k+1)}^{j}\mid s_{t,(k)})$.
Definition 11.
The characteristic peak-based function is defined as:
$\nu_{p}(\mathcal{C}_{\sigma}^{i},s_{t,(k)})\coloneqq\sum_{j\in\mathcal{C}_{\sigma}^{i}}\mathrm{H}\left(A_{t,(k+1)}^{j}\right)$
Consensus-based.
While the policy distribution may have the same shape even if the agent chooses a different action, $H(A_{t})$ does not capture such changes. To address this, and to assess whether agents act in line with expectations to others, we use the Jensen–Shannon Divergence (JSD), which has been used to encourage diversity among agents [25]. For two PMFs $p$ and $q$ with $u=\frac{1}{2}(p+q)$, the JSD is a symmetrized version of $D_{KL}$:
$\mathrm{JSD}(p\mid\mid q)=\frac{1}{2}D_{KL}(p\mid\mid u)+\frac{1}{2}D_{KL}(q\mid\mid u).$ (7)
While in the ordinary MG framework (Def. 1) each agent $i$ observes the global state $s_{t}$, in practice agents either learn which components of $S_{t}$ pertain to them, or the environment assigns agent $i$’s components to a fixed (typically first) position. We denote by $s_{t}^{i}$ the value of agent $i$’s component in the global state, and by $s_{t}^{(i)}$ the agent’s individualized observation. Based on this, we define a consensus metric to include the cases of: i) others-consensus, to infer if agent $i$ would act similarly in the state of others and, ii) self-consensus, to infer if others would act similarly in the state agent $i$.
Using base-2 $\log$, $\mathrm{JSD}(p\|q)$ is bounded in $[0,1]$, where 0 indicates identical distributions and $1$ maximal divergence. To instead express similarity, we define $J(p\|q)\coloneqq 1-\mathrm{JSD}(p\|q)$.
Definition 12.
The characteristic consensus-based function is:
$\nu_{c}(\mathcal{C}_{\sigma}^{i},s_{t,(k)})\coloneqq$ $\sum_{j\in\mathcal{C}_{\sigma}^{i}}\bigg{[}\overbrace{J\left(\pi^{i}\left(\cdot\mid s^{(j)}_{t,(k)}\right)\mid\mid\pi^{j}\left(\cdot\mid s^{(j)}_{t,(k)}\right)\right)}^{other}$
$+\underbrace{J\left(\pi^{i}\left(\cdot\mid s^{(i)}_{t,(k)}\right)\mid\mid\pi^{j}\left(\cdot\mid s^{(i)}_{t,(k)}\right)\right)}_{self}\bigg{]}.$
If only one of the consensus terms is used, we denote others-consensus by $\nu_{co}$ and self-consensus by $\nu_{cs}$. For dissimilarity, we use $\mathrm{JSD}$ in place of $J$ in $\nu_{c}$ and term it $\nu_{d}$, $\nu_{do}$ and $\nu_{ds}$ respectively.
4.4 Instrumental Empowerment
For each definition of $\nu$, we link the marginal contribution from Def. 8 to important and well-known metrics in MARL. In the following, we use the shorthand $\Delta\nu$ to denote $\Delta\nu(\mathcal{C}_{\sigma}^{i},s_{t,(k)})$.
Definition 13.
The advantage of agent $j$ is defined by $\boldsymbol{\Lambda}^{j}(s_{t},\boldsymbol{a}_{t})\coloneqq Q^{j}(s_{t},\boldsymbol{a}_{t})-V^{j}(s_{t})=\mathbb{E}_{s_{t+1}\sim P}[R^{j}+\gamma V^{j}(s_{t+1})]-V^{j}(s_{t})$.
Proposition 1.
For each intermediate state $s_{t,(k)}$, $\Delta\nu_{v}$ measures the sampled advantage for players in $\mathcal{C}_{\sigma}^{i}$ attributable to action $a_{t}^{i}$:
$\Delta\nu_{v}=\sum_{j\in\mathcal{C}_{\sigma}^{i}}V^{j}(s_{t,(k)})-V^{j}(s_{t,(k-1)})=\sum_{j\in\mathcal{C}_{\sigma}^{i}}\Lambda^{j}(s_{t,(k)},a_{t}^{i})$
Proof.
See Appendix C.1. ∎
Definition 14.
The empowerment function [16] quantifies an agent’s maximal potential causal influence on future states, defined as $\mathcal{E}(S_{t})=\max_{p(a_{t}\mid s_{t})}\left[H(S_{t+1}\mid S_{t})-H(S_{t+1}\mid S_{t},A_{t})\right]$.
Proposition 2.
For each intermediate state $s_{t,(k)}$, $\Delta\nu_{p}$ quantifies the instrumental empowerment as the increase in decision certainty of players in $\mathcal{C}_{\sigma}^{i}$ attributable to action $a_{t}^{i}$:
$\Delta\nu_{p}=\sum_{j\in\mathcal{C}_{\sigma}^{i}}\mathrm{H}\left(A_{t,(k+1)}^{j}\right)-\mathrm{H}\left(A_{t,(k)}^{j}\right)=\sum_{j\in\mathcal{C}_{\sigma}^{i}}\mathcal{E}^{j}(s_{t,(k-1)})$
Proof.
See Appendix C.2. ∎
Definition 15.
An agent $i$’s action $a^{i*}$ is a best response to actions of co-players $\boldsymbol{i}$ if it maximizes its payoff $u^{i}$ given the strategies chosen by the others, formally defined as $a^{i*}=\mathrm{argmax}_{a^{i}\in\mathcal{A}^{i}}u^{i}(a^{i},a^{-\boldsymbol{i}})$.
Proposition 3.
At state $s_{t,(k)}$, the action $a_{t}^{i}$ increases consensus to co-players $j\in\mathcal{C}_{\sigma}^{i}$ if $\Delta\nu_{j}\geq 0$. If furthermore $\Delta\nu_{c}(\cdot\mid a_{t}^{i})=\max_{a\in\mathcal{A}^{i}}\Delta\nu_{c}(\cdot\mid a)$, then $a_{t}^{i}$ is a best response at $s_{t,(k)}$ in terms of alignment with co-player strategies.
Proof.
Follows directly from Def. 15, as maximizing $\nu_{c}$ is equivalent to maximizing $\Delta\nu_{c}$. ∎
5 Experiments
To assess whether the intuition from Fig. 1 and our ICV method generalize to more complex settings, we conduct empirical validation in three environments. [^1] Python code available at github.com/icv-marl. Each illustrates distinct applications of ICVs online and offline for extracting behavioral insights and identifying credit assignment in the absence of rewards. We examine whether $\Phi(\nu_{p})$ and $\Phi(\nu_{c})$ can serve as proxies for $\Phi(\nu_{v})$. As the validity of using the value-based function $\nu_{v}$ in $\Phi(\nu_{v})$ as a feature contribution predictor has been proven in [3], it serves as our baseline. We compute ICVs over $M$ episodes and report relative ICV values by $i)$ normalizing all characteristic functions $\nu$, and ii)ii) dividing each $\Phi(\cdot)$ by $\kappa=\max\Phi(\cdot)$ to scale results close to $1$, as some contributions may cancel out and yield low contributions on average. We assume convergence of the learning process, i.e., no further performance improvement, but note that our method is applicable at any point during training.
5.1 Cooperative Game
For the fully cooperative case, we consider the environment Level-based Foraging (LBF), illustrated in Fig. 4(a). Agents and apples are assigned levels, and picking up an apple requires the sum of the levels of nearby agents to match the apple’s level, followed by a simultaneous load action. Each agent uses an individual centralized critic, and our ICV method is applied online. We first validate the correspondence between value $V^{i}$ and peak $\mathrm{H}^{i}$ increase in Fig. 4. In the depicted initial state (left), agent 1 moves to the right (green arrow), thereby increasing both its own and its teammates’ certainty in action selection, while also increasing the value for all, which is shown in the right plot. This effect arises because agent 1 positions itself below the apple, therefore clarifying to others where they should move to.
(a) LBF environment.
up   down   left   right   load   value
$\pi^{0}(s_{t})$$\pi^{1}(s_{t})$$\pi^{2}(s_{t})$011
$\pi^{0}(s_{t})$$\pi^{1}(s_{t})$$\pi^{2}(s_{t})$011
(b) Policy distributions and actions.
Figure 4: The effect of agent 1’s action (moving right) on the policies and values before (top) and after (bottom) acting.
Contrary to the previous, we verify the correspondence between increase in value $V^{i}(\cdot)$ and uncertainty in decision-making, measured by $H(A_{t}^{i})$ for each agent. Besides the entropy, we count the number of actions $a_{t}^{i}$ yielding $\Lambda^{i}(s_{t},\boldsymbol{a}_{t}^{\boldsymbol{-i}},a_{t}^{i})\geq 0$ as per Def. 13, by keeping the actions $\boldsymbol{a^{-i}}$ fixed and term this as agent $i$’s choice $C^{i}$. We plot the normalized values of $V^{i}$, entropy $H^{i}=H(A_{t}^{i})$ and choices $C^{i}$ in Fig. 5. The plots per agent $i$ indicate a reasonable correspondence between increases in value and decrease of entropy and number of valuable choices, thereby further supporting our intuition.
$V^{i}$$H^{i}$$C^{i}$
05510101515011
(a) Agent 0.
05510101515011
(b) Agent 1.
05510101515011
(c) Agent 2.
Figure 5: Comparison between normalized values $V^{i}$, entropy $H^{i}$ and choices $C^{i}$ along y-axis over time steps $t$ in x-axis of LBF agents.
[…]
A0A1A21-10.5-0.500.50.511
$\hat{\Phi}(\nu_{v})$$\hat{\Phi}(\nu_{p})$$C^{i}$$\hat{\Phi}(\nu_{cs})$$\hat{\Phi}(\nu_{co})$
[…]
5.2 Mixed-Motive Game
For the mixed-motive setting, we use the Multi-Particle Environment (MPE) Tag, shown in Fig. 7(a), where predators (red) aim to catch a prey (green). Agents have individual centralized critics, and ICV computation is performed offline, although online computation would be feasible since each agent is processed independently. In Fig. 7(b), we analyze agent 1 (prey) and again verify the correspondence between its value $V^{1}$ and action determinism $\mathrm{H}^{1}$ over a single episode. The results indicate a reasonable alignment between value and action certainty. We further assess agent 1’s dissimilarity to its opponents (predators) by setting $\mathcal{C}_{\sigma}^{i}$ accordingly and measuring $\nu_{d}$, denoted as $\mathrm{J}^{1}$. Agent 1 shows high dissimilarity throughout the episode which is consistent with the intuition that a prey should act contrary to predators, ideally avoiding movement toward them.
(a) MPE environment.
$V^{1}$   $\mathrm{H}^{1}$   $\mathrm{J}^{1}$
0101020203030404050501-1011
(b) Episode history of agent 1.
Figure 7: Comparing normalized value $V^{1}$, $\mathrm{H}^{1}$ and $\mathrm{J}^{1}$ on the right plot along y-axis over time steps $t$ along x-axis of an MPE prey.
We measured the ICVs over $M=500$ episodes and normalized them using $\kappa=0.08$. Due to the symmetry of the game, we focus on the contributions of a single adversary (predator, denoted as Adv) and a single agent (prey, denoted as Ag). For both, we partition their coalitions and evaluate the effects on teammates (-t) and opponents (-o), where $\nu_{c}$ is used for teammates, while dissimilarity $\nu_{d}$ is applied to opponents. For predators, this means: in Adv-t, $\nu_{c}$ captures similarity among predators and $\nu_{d}$ measures dissimilarity between predators and prey. In Adv-o, $\nu_{c}$ reflects similarity among prey, and $\nu_{d}$ their dissimilarity with predators. The same interpretation applies to the contributions of prey (Ag).
Figure 8 presents the results, where proximity to others is the main effect. The prey’s contributions align with our expectations: average value effects $\hat{\Phi}(\nu_{v})$ correspond to action determinism $\hat{\Phi}(\nu_{p})$ and alignment with co-player’s preferences (Ag-t, $\hat{\Phi}(\nu_{c})$), but interestingly, they contribute to less diversity toward predators $\hat{\Phi}(\nu_{d})$. In contrast, predators benefit from reducing action determinism and keeping their choices open when engaging prey, as indicated by increased $\hat{\Phi}(\nu_{c})$ but decreased $\hat{\Phi}(\nu_{p})$. They also tend to behave slightly more similar to each other (Adv-t, $\hat{\Phi}(\nu_{c})$) while promoting diversity against opponents ($\hat{\Phi}(\nu_{d})$). In Adv-o, high $\hat{\Phi}(\nu_{p})$ and low $\hat{\Phi}(\nu_{v})$ indicate prey become more deterministic (e.g., fleeing) as predators approach. Their effect also renders prey reacting differently (Adv-o, $\hat{\Phi}(\nu_{c})$). Overall, compared to Fig. 7(b), the ICVs suggest that $\hat{\Phi}(\nu_{p})$ can signal both positive and negative contributions.
Adv-tAdv-oAg-tAg-o1-10.5-0.500.50.511
$\hat{\Phi}(\nu_{v})$$\hat{\Phi}(\nu_{p})$$\hat{\Phi}(\nu_{c})$$\hat{\Phi}(\nu_{d})$
[…]
5.3 Competition Game
In the competitive setting, we analyze agent behavior in Google Research Football (GRF, Fig. 9[19], a complex game simulating real-world football. We employ TiZero [21], a strong model, where the policy $\pi_{tz}$ is shared across agents and differentiated only by the agent index, effectively inducing heterogeneous roles. A shared centralized critic is used, and ICV computation is performed offline.
(a) Opponent scoring goal.
(b) Team scoring goal.
Figure 9: Two goal scenarios in GRF with ball marked in red.
Our approach is flexible to measure the effects of specific groups. We define a fixed agent order $\sigma=(\boldsymbol{p}_{o},\boldsymbol{p}_{t},p_{s})$, where $\boldsymbol{p}_{o}$ are opponent players, $\boldsymbol{p}_{t}$ are teammates, and $p_{s}$ is the striker or the ball player (when the team has possession). Groups $\boldsymbol{p}_{o}$ and $\boldsymbol{p}_{t}$ are processed jointly, and ICV computation is restricted to every second joint action ($\Delta t=2$, see Fig. 2). Since $\pi_{tz}$ is shared across agents, we measure $\bar{\nu}_{d}=\mathrm{JSD}(\pi_{tz}(\cdot\mid s_{t,(k+1)})\,\|\,\pi_{tz}(\cdot\mid s_{t,(k)}))$ as a proxy for strategy change, alongside entropy $\mathrm{H}$ to capture decision uncertainty. The contributions over virtual game time $\tilde{t}$ (in minutes) with seed=175 are shown in Fig. 10. We highlight three key regions, where the situation can be reconstructed to investigate the cause as GRF can store full history. In the gray area, the team loses and regains possession near the penalty area, causing uncertainties and strategy adaption. In the orange region, the opponents scored a goal as reflected by a drop in $\hat{\Phi}(\nu_{v})$ and a rise in $\hat{\Phi}(\bar{\nu}_{d})$, indicating a strong strategy adaptation. In the green region, opponents exert a negative impact on the ball player’s value and determinism just before it scored a goal. In contrast, teammates contribute positively by increasing the ball player’s decisiveness to shoot. While $\hat{\Phi}(\nu_{v})$ and $\hat{\Phi}(\nu_{p})$ contributions do not strictly match, $\hat{\Phi}(\nu_{p})$ and $\hat{\Phi}(\bar{\nu_{d}})$ may still provide meaningful information on the contributions of the players.
$\hat{\Phi}(\nu_{v})$   $\hat{\Phi}(\nu_{p})$   $\hat{\Phi}(\bar{\nu}_{d})$
055101015151-1011
(a) Effect of opponents on ball player or striker.
05510101515011
(b) Effect of team players on ball player or striker.
Figure 10: GRF history effects on ball player or striker.
Given the long GRF time horizon ($T=3000$), we compute ICV values over $M=50$ episodes and normalize them by $\kappa=0.04$. We analyze the contributions of opponents (opp) and teammates to the attacking trio, i.e., right midfield (rm), center forward (cf), and left midfield (lf), as shown in Fig. 11(a), and to the ball player (bp) only in Fig. 11(b). We decompose the peak contributions $\hat{\Phi}(\nu_{p})$ to the trio by $\mathrm{H}^{rm}$, $\mathrm{H}^{cf}$, and $\mathrm{H}^{lm}$. We observe that credit got assigned to contributions increasing or decreasing decision certainty that again correlates to the effects on the value. Interestingly, neither opponents nor co-players had a large impact on the ball player on average, making it robust to disturbances by frequently adapting its strategy, as indicated by high $\hat{\Phi}(\bar{\nu}_{d})$.
$\hat{\Phi}(\nu_{v})$   $\mathrm{H}^{rm}$   $\mathrm{H}^{cf}$   $\mathrm{H}^{lm}$
OppTeam1-10.5-0.500.50.511
(a) Effect on forwarding players.
$\hat{\Phi}(\nu_{v})$   $\hat{\Phi}(\nu_{p})$   $\hat{\Phi}(\bar{\nu}_{d})$
OppTeam1-10.5-0.500.50.511
(b) Effect on ball player.
Figure 11: ICVs in GRF of opponents and team players on forwarding and ball player in terms of value, determinism and strategy change.
6 Conclusions
This paper introduced Intended Cooperation Values (ICVs), a novel causal-action attribution method that explains agent behavior through policy influence rather than reward-based evaluations. ICVs quantify how individual actions affect teammates’ decision certainty and preference alignment, effectively capturing implicit cooperation and consensus. The approach proved effective across a variety of environments (cooperative, competitive, and mixed-motive). A major advantage of ICV lies in its simplicity and generality to operate without requiring reward access or architectural changes and can be applied both online and offline. This makes them well-suited for scenarios where interpretability is critical but value functions are unavailable or unreliable. Furthermore, ICVs offer a principled way to analyze intention and influence, grounded in game-theoretic and information-theoretic principles. However, the method also has limitations. Computing ICVs in complex, large-scale environments requires considerable offline post-processing due to the need for sequential action decomposition and intermediate state reconstruction. In addition, our method assumed full state observability, which may limit applicability in some settings. Nonetheless, our method is conceptually extendable to partially observable environments.
Supplementary Material
Appendix A Ethics Statement.
Our method does not incorporate any form of adversarial attack and pose no risk to human safety. All experiments were conducted in a simulated environment, without involving real-world subjects or raising ethical or fairness concerns. The method is intended to enhance transparency and safety in MARL systems, especially in critical applications.
Appendix B Shapley Value Axioms
  1. 1.
    Efficiency: The sum of the SVs for all players equals the total value of the grand coalition:
    $\sum_{i=1}^{n}\phi_{i}(v)=v(N)\,.$ (8)
  2. 2.
    Symmetry: If two players are interchangeable (i.e., they contribute equally to every possible coalition),
    $v(C\cup\{i\})=v(C\cup\{j\})\quad\text{for all }C\subseteq N\setminus\{i,j\}\,$, (9)
    then they receive the same SV $\phi_{i}(v)=\phi_{j}(v)$.
  3. 3.
    Additivity: For any two games $v$ and $w$, the SV of their sum $v+w$ is the sum of their SVs:
    $\phi_{i}(v+w)=\phi_{i}(v)+\phi_{i}(w)\quad\text{for all }i\,.$ (10)
  4. 4.
    Dummy Player: If a player does not contribute any additional value to any coalition $C\subseteq N$,
    $v(C\cup\{i\})=v(C)\,$, (11)
    then its SV is zero $\phi_{i}(v)=0$.
Appendix C Proofs
C.1 Advantage
Proposition 1.
For each intermediate state $s_{t,(k)}$, $\Delta\nu_{v}$ measures the sampled advantage for players in $\mathcal{C}_{\sigma}^{i}$ attributable to agent $i$’s action $a_{t}^{i}$:
$\Delta\nu_{v}=\sum_{j\in\mathcal{C}_{\sigma}^{i}}V^{j}(s_{t,(k)})-V^{j}(s_{t,(k-1)})=\sum_{j\in\mathcal{C}_{\sigma}^{i}}\Lambda^{j}(s_{t,(k)},a_{t}^{i}).$
Proof.
Given $Q^{j}(s,\boldsymbol{a})=\mathbb{E}_{\boldsymbol{\pi},P}[\sum_{t=0}^{\infty}\gamma^{t}R^{j}\mid S_{0}=s,\boldsymbol{A}_{0}=\boldsymbol{a}]$, the advantage for each agent $j$ is:
$\boldsymbol{\Lambda}^{j}(s_{t,(k)},\boldsymbol{a}_{t})=Q^{j}(s_{t,(k)},\boldsymbol{a}_{t})-V^{j}(s_{t,(k)})=\mathbb{E}_{s_{t+1}\sim P}[R^{j}+\gamma V^{j}(s_{t+1})]-V^{j}(s_{t})$
by noting that our SVMG framework operates between intermediate states $s_{t,(k)}$ at sub-steps $k$, where the time step $t$ is not increased, and noting that rewards are provided only at full next time steps $t+1$, we may omit $R^{j}$ and $\gamma$. Additionally, according to Def. 6 and 8, we consider concrete next intermediate states after agent $i$’s action $a_{t}^{i}$ got processed. We therefore obtain the following advantage for each agent $j$ based on agent $i$’s action (or state propagation):
$\Lambda^{j}(s_{t,(k)},a_{t}^{i})=V^{j}(s_{t,(k+1)})-V^{j}(s_{t,(k)}).$
Then, in the case of centralized critics, we obtain the same result through the Advantage Decomposition Theorem [18], being applicable for all agents at sub-step $k$ for simplicity we use an ordered permutation $k=\sigma(k)$:
$\boldsymbol{\Lambda}^{1:n}(s_{t,(k)},\boldsymbol{a}_{t}^{1:n})$ $=\sum_{k=1}^{n}\Lambda^{k}(s_{t,(k)},\boldsymbol{a}_{t}^{1:k-1},a_{t}^{k})$ (12)
$=\sum_{k=1}^{n}Q^{1:k}(s_{t,(k)},\boldsymbol{a}^{1:k})-Q^{1:k-1}(s_{t,(k)},\boldsymbol{a}^{1:k-1})$ (13)
where $Q^{1:k}(s_{t,(k)},\boldsymbol{a}^{1:k})=\mathbb{E}_{a^{k+1:n}\sim\pi^{k+1:n}}\left[Q(s_{t,(k)},\boldsymbol{a}^{1:k},\boldsymbol{a}^{k+1:n})\right]$. With the same reasoning as before, $Q^{1:k}(s_{t,(k)},\boldsymbol{a}^{1:k})=V(s_{t,(k+1)})$ and $Q^{1:k-1}(s_{t,(k)},\boldsymbol{a}^{1:k-1})=V(s_{t,(k)})$ yields the advantage at sub-step $k$:
$\Lambda^{k}(s_{t,(k)},a_{t}^{1:k-1},a_{t}^{k})=V(s_{t,(k+1)})-V(s_{t,(k)}).$
C.2 Empowerment
Proposition 2.
For each intermediate state $s_{t,(k)}$, $\Delta\nu_{p}$ measures the instrumental empowerment as the certainty increase in decision-making of players in $\mathcal{C}_{\sigma}^{i}$ attributable to agent $i$’s action $a_{t}^{i}$:
$\Delta\nu_{p}=\sum_{j\in\mathcal{C}_{\sigma}^{i}}\mathrm{H}\left(A_{t,(k+1)}^{j}\right)-\mathrm{H}\left(A_{t,(k)}^{j}\right)=\sum_{j\in\mathcal{C}_{\sigma}^{i}}\mathcal{E}^{j}(s_{t,(k-1)}).$
Proof.
We assume $\pi^{j}(\cdot\mid s_{t,(k)},a_{t}^{i})\approx\pi^{j}(\cdot\mid s_{t,(k+1)})$ as otherwise communication-based policies would be required, which are not the scope of this work. At $s_{t,(k-1)}$, by observing action $a_{t}^{i}$ but not executing it to get a new state as per Def. 6, it holds that
$\Delta\nu_{p}$ $=\sum_{j\in\mathcal{C}_{\sigma}^{i}}\log|\mathcal{A}^{j}|-H\left(A_{t,(k)}^{j}\mid s_{t,(k-1)},a_{t}^{i}\right)-\log|\mathcal{A}^{j}|+H\left(A_{t,(k)}^{j}\mid s_{t,(k-1)}\right)$
$=\sum_{j\in\mathcal{C}_{\sigma}^{i}}H\left(A_{t,(k)}^{j}\mid s_{t,(k-1)}\right)-H\left(A_{t,(k)}^{j}\mid s_{t,(k-1)},a_{t}^{i}\right)$
$=\sum_{j\in\mathcal{C}_{\sigma}^{i}}\mathrm{I}\left(A_{t,(k)}^{j};a_{t}^{i}\mid s_{t,(k-1)}\right).$ (14)
By setting the probing distribution $p=\pi$, we obtain $\max_{\pi^{i}}\mathrm{I}\left(A_{t,(k)}^{j};a_{t}^{i}\mid s_{t,(k-1)}\right)=\mathcal{E}(s_{t,(k-1)})$. ∎